OpenStudy (anonymous):

the water depth in a harber is 21m at hight tide, and 11m at low tide. One cycle is completed approximatly every 12h. a) find an equation for the water depth as a function of the time, t hours, after low tide

OpenStudy (anonymous):

i found the amplitude and the period of the equation y=asink(x-c)+d Amplitude a=21-11/2 a=5 K p=2pi/k 12=2pi/k k=2pi/12 k=pi/6 what about the c and the d??

OpenStudy (anonymous):

@lgbasallote @satellite73 @dpaInc help? :S

OpenStudy (anonymous):

the answer at the back of the book says y=5sin pi/6 (t-3)+16